Please shed some light on finding the solution.
If Emily hits a golf ball from the upper deck of a driving range (twelve feet above the ground) with an initial velocitynof 117 feet per second, after how many seconds should the ball land (assume her ball does not encounter any obstacles)? Round to the nearest tenth.

Answer by Alan3354(69443)

If Emily hits a golf ball from the upper deck of a driving range (twelve feet above the ground) with an initial velocitynof 117 feet per second, after how many seconds should the ball land (assume her ball does not encounter any obstacles)? Round to the nearest tenth.
------------
If it's path is horizontal, it's the time it takes it to fall 12 feet.
h(t) = -16t^2 + 12 = 0
t^2 = 3/4
t = sqrt(3)/2 seconds
t =~ 0.866 seconds
----------------
If it's some other angle, the angle has to be given.